# Boolean Satisfiability via Imitation Learning

### v1

@@ 1 Introduction
The Boolean satisfiability (SAT) problem is a cornerstone of theoretical computer science and artificial intelligence (Cook, 1971; Karp, 1972). Beyond its foundational role, SAT serves as the computational backbone of numerous applications, including formal verification, planning, and combinatorial optimization. Modern solvers for SAT are dominated by the conflict-driven clause learning (CDCL) framework (Silva & Sakallah, 1996; Biere et al., 2009), which has scaled to industrial benchmarks of immense complexity. A CDCL run interleaves branching, unit propagation, and conflict analysis. Among these components, the branching rule largely determines the search trajectory, while unit propagation often dominates runtime (Zhang & Malik, 2002; Davis et al., 2008; Moskewicz et al., 2001). As a result, more informed branching decisions can translate directly into faster solving.
@@ 2 Related Works
Early learning approaches focused on instance-level prediction, using Graph Neural Network (GNNs) (Scarselli et al., 2008) to classify SAT or UNSAT, as seen in NeuroSAT (Selsam et al., 2019; Selsam & Bjørner, 2019) and  (Cameron et al., 2020). Recent work has explored whether Transformers can learn solver behavior directly (Pan et al., 2025). In parallel with these model-based approaches, complementary efforts target data and benchmarking, including G2SAT (You et al., 2019) and G4SATBench (Li et al., 2024). Building on these foundations, a second line integrates learning inside solvers to shape specific components: for example, NeuroSelect (Liu et al., 2024) learns clause deletion policies, NeuroBack (Wang et al., 2024) improves phase initialization with GNNs, and RDC‑SAT (Zhai & Ge, 2025) adopts a divide-and-conquer strategy via reinforcement learning. This approach leads to targeted enhancements within solver mechanisms. More concretely, within the CDCL branching loop, several methods exemplify this integration: NeuroSAT (Selsam & Bjørner, 2019) has been used to guide variable selection; Graph-Q-SAT (Kurin et al., 2020) trains an RL agent queried online during search based on instance information; and SATformer (Shi et al., 2023) trains a GNN Transformer model to initialize the CDCL that indirectly influences branching thereafter.
@@
Imitation learning (IL) learns policies directly from expert demonstrations, that is, sequences of states with associated actions (Osa et al., 2018; Zare et al., 2024). A simple example is behavior cloning (BC), which utilizes supervised learning to map observed situations to expert choices (Pomerleau, 1991). Building on the principles of imitation learning, the Decision Transformer (Chen et al., 2021) is similar to behavior cloning, framing reinforcement learning as sequence modeling, where an autoregressive Transformer is trained to predict the next action given a sequence rollout of returns, states, and actions. This view connects control to next-token prediction and attains competitive performance without explicit value function learning. The application of imitation learning extends beyond traditional domains. For instance, beyond robotics and games, IL has guided decision-making in exact optimization solvers. In mixed-integer linear programming, policies learned to imitate strong branching can be used within branch-and-bound and achieve strong results (Gasse et al., 2019). Related work also learns branching policies that integrate into branch-and-bound (Zarpellon et al., 2021).
@@ 3 Preliminaries on SAT
and a formula is in conjunctive normal form (CNF) (Plaisted & Greenbaum, 1986) if it is a conjunction of clauses:
@@
In practice, SAT formulas are often represented in the standard DIMACS CNF format (Johnson & Trick, 1996). In this encoding, each clause is written as a sequence of nonzero integers terminated by 0, where the integer $i$ denotes variable $x_{i}$ and i-i denotes its negation $\lnot x_{i}$. For the formula in Equation 3, the DIMACS form is:
@@
Conflict-driven clause learning (CDCL) is the dominant algorithmic framework for practical SAT solvers (Silva & Sakallah, 1996; Biere et al., 2009). A CDCL solver incrementally explores a branching search tree through three fundamental operations: decision ($\mathtt{D}$), where the solver assigns a value to a chosen literal and thus extends the current partial assignment; unit propagation ($\mathtt{A}$), also known as Boolean Constraint Propagation (BCP) (Moskewicz et al., 2001), where implied literals are deduced from unit clauses; and backtracking ($\mathtt{BT}$), where a detected conflict triggers clause learning and a non-chronological jump to an earlier decision level.
@@
Usually, the choice of the next decision variable, together with its phase, is determined by a branching heuristic. The most influential family of heuristics is the Variable State Independent Decaying Sum (VSIDS) (Moskewicz et al., 2001). In VSIDS, variables are assigned activity scores that are increased whenever they appear in learned clauses. These scores are then periodically decayed, ensuring that recently relevant variables are favored for branching. This dynamic prioritization enables the solver to focus on the most promising parts of the search space. A notable implementation is MiniSAT (Eén & Sörensson, 2003), a lightweight and extensible open-source CDCL solver. Despite its minimalistic codebase, MiniSAT has become a standard platform for both research and industrial applications due to its clarity and effectiveness.
@@ 4.1 Expert KeyTrace Replay
We seek an expert that guides the solver toward effective branching decisions while avoiding wasted detours. Raw CDCL trails are often long and contain many decisions that are later undone by backtracking. An ideal expert should preserve only the decisions that survive and discard detours, thereby yielding clean training targets and eliminating steps that do not advance the search. In CDCL, unit propagation dominates runtime, typically accounting for 80%–90% of the total solving time (Zhang & Malik, 2002; Davis et al., 2008; Moskewicz et al., 2001), a trend we also confirm in Figure 2. This observation suggests that an expert which avoids wasted branches can substantially reduce propagation and thereby improve overall efficiency.
@@ 5.1 Implementation Details
We compare against SATformer (Shi et al., 2023) and Graph‑Q‑SAT (Kurin et al., 2020), which are trained with their public implementations and the training data specified in their papers.
@@
A Python reimplementation of MiniSAT 2.2 (Eén & Sörensson, 2003) is used and validated against the official C++ version on full trails $\mathcal{T}_{t}$. This version streamlines the integration of learning methods, ensuring that all approaches are evaluated in a consistent environment.
@@
We generate the random 3–SAT training dataset with a planted assignment (Mezard & Montanari, 2009; Achlioptas et al., 2000; 2005). The clause–variable ratio is in $[4.1,4.4]$. Variable ranges include 5 to 15, 16 to 30, 31 to 60, and 61 to 100, with additional fixed sizes of 50 and 100. For each formula, a MiniSAT run yields the level-annotated trail as Equation 5. Then, the expert KeyTrace is extracted using Equation 7. Every decision position provides one supervision pair, a KeyTrace prefix $\mathcal{K}_{t}$, and its next decision $d_{t+1}$. For evaluation, we first use the test set, which is drawn from the same generators and variable ranges as those used for training. Generalization is then assessed on different SAT families from SATLIB (Hoos & Stützle, 2000). Further dataset details appear in Appendix B.
@@
Propagation dominates CDCL time (Zhang & Malik, 2002; Davis et al., 2008; Moskewicz et al., 2001), as seen in Figure 2 and Table 6. We therefore use the number of propagation as the primary signal. Since instances vary in size and difficulty, we normalize by MiniSAT on a per‑instance basis and aggregate with the median to reduce the influence of outliers.
@@ 5.2 Comparison on test sets
We compare ImitSAT with Graph-Q-SAT and SATformer on held-out random 3-SAT. To align wall-clock budgets, ImitSAT is queried at 3 or 5 times per instance. Graph-Q-SAT uses the same computational budgets. SATformer updates the VSIDS variable scores once at initialization. We report MRPP $\tilde{r}$ and the one percent win rate $W_{1\%}$.
@@
MRPP $\tilde{r}$ ($\downarrow$) GQSAT-3calls 1.00 0.94 0.89 1.15 0.71 0.85
@@
$W_{1\%}$ ($\uparrow$) GQSAT-3calls 0.45 0.53 0.54 0.48 0.57 0.55
@@ 5.3 Generalization on special SAT families
Across all families, ImitSAT attains the lowest MRPP $\tilde{r}$ or ties for best as shown in Table 2 (top). The gains are large on PARITY and PRET, and with 5 calls ImitSAT is best on JNH and AIM, and matches the best on PHOLE. Graph-Q-SAT does not reduce propagation number on JNH and AIM, and in fact increases the propagation number. SATformer only shows a gain on the PARITY dataset; all other datasets exhibit no gain or worse performance than MiniSAT. Table 2 (bottom) reports the $W_{1\%}$. ImitSAT achieves the highest or tied win rate on all families, including a perfect score on PRET and matching the best on PARITY and PHOLE. These outcomes indicate that imitation of the expert KeyTrace transfers from random 3-SAT to structured regimes without finetuning.
@@
MRPP $\tilde{r}$ ($\downarrow$) GQSAT-3calls 1.29 1.20 0.82 1.03 0.88
@@
$W_{1\%}$ ($\uparrow$) GQSAT-3calls 0.38 0.31 0.80 0.50 0.50
@@ 5.4 Wall-clock Time
Across random 3‑SAT test sets and structured families, ImitSAT traces the lowest curves under these budgets, which means more instances are solved in less time, as shown in Figure 4. Learning model-based branching introduces query overhead, so wall‑clock gains appear only once propagation savings exceed this cost.
Figure 4: Wall‑clock time versus instances solved on random test sets and structured families. Curves show end‑to‑end solver time per instance with all model calls included and preprocessing excluded. ImitSAT and Graph‑Q‑SAT use 3 calls per instance. SATformer performs a single VSIDS initialization. The lower curve is better.
@@ Appendix B Datasets
This section details the training corpus, held‑out test sets, and the structured SAT families used in evaluation. For the synthetic data, both train and test, all instances are random 3-SAT with a planted assignment and a clause–variable ratio in [4.1, 4.4][4.1,\,4.4] (Mezard & Montanari, 2009; Achlioptas et al., 2000; 2005). We categorize instances by the number of variables $n_{v}$ to control difficulty and balance the number of decision probes contributed by each category. Since larger $n_{v}$ produces longer trails and thus more supervision per instance, we allocate fewer instances at larger $n_{v}$. This ensures each bucket contributes a similar amount of training signal. We define buckets as ranges of variable counts (e.g., “5–15”), with each bucket grouping instances by the number of variables $n_{v}$ they contain.
@@
To assess cross‑distribution generalization, we evaluate on classic structured benchmarks from SATLIB (Hoos & Stützle, 2000). The suite spans both SAT and UNSAT, and includes 3-SAT and non–kk–SAT regimes. Specifically, we use AIM (Asahiro et al., 1996), which is equivalent to 3-SAT, focusing on the $n_{v}=100$ SAT portion. We also consider the SAT subset of JNH (Selman et al., 1996) with $n_{v}=100$ and PARITY (Warners & Van Maaren, 1998; Hoos & Stützle, 2000) in the compressed series, both as structured SAT families distinct from our random generator. Additionally, PRET (Johnson & Trick, 1996; Hoos & Stützle, 2000) is a two-coloring UNSAT problem with $n_{v}=60$, and the pigeonhole principle (PHOLE) (Haken, 1985) is used as a non–kk–SAT and UNSAT stress test. All subsets described above are considered with $n_{v}\leq 100$. Table 5 lists the families.

### v2

@@ 1 Introduction
The Boolean satisfiability (SAT) problem is a cornerstone of theoretical computer science and artificial intelligence (Cook, 1971; Karp, 1972). Beyond its foundational role, SAT serves as the computational backbone of numerous applications, including formal verification, planning, and combinatorial optimization. Modern solvers for SAT are dominated by the conflict-driven clause learning (CDCL) framework (Silva and Sakallah, 1996; Biere et al., 2009), which has scaled to industrial benchmarks of immense complexity. A CDCL run interleaves branching, unit propagation, and conflict analysis. Among these components, the branching rule largely determines the search trajectory, while unit propagation often dominates runtime (Zhang and Malik, 2002; Davis et al., 2008; Moskewicz et al., 2001). As a result, more informed branching decisions can translate directly into faster solving.
@@ 2 Related Works
Early learning approaches focused on instance-level prediction, using Graph Neural Network (GNNs) (Scarselli et al., 2008) to classify SAT or UNSAT, as seen in NeuroSAT (Selsam et al., 2019; Selsam and Bjørner, 2019) and  (Cameron et al., 2020). Recent work has explored whether Transformers can learn solver behavior directly (Pan et al., 2025). In parallel with these model-based approaches, complementary efforts target data and benchmarking, including G2SAT (You et al., 2019) and G4SATBench (Li et al., 2024). Building on these foundations, a second line integrates learning inside solvers to shape specific components: for example, NeuroSelect (Liu et al., 2024) learns clause deletion policies, NeuroBack (Wang et al., 2024) improves phase initialization with GNNs, and RDC‑SAT (Zhai and Ge, 2025) adopts a divide-and-conquer strategy via reinforcement learning. This approach leads to targeted enhancements within solver mechanisms. More concretely, within the CDCL branching loop, several methods exemplify this integration: NeuroSAT (Selsam and Bjørner, 2019) has been used to guide variable selection; Graph-Q-SAT (Kurin et al., 2020) trains an RL agent queried online during search based on instance information; and SATformer (Shi et al., 2023) trains a GNN Transformer model to initialize the CDCL that indirectly influences branching thereafter.
@@
Imitation learning (IL) learns policies directly from expert demonstrations, that is, sequences of states with associated actions (Osa et al., 2018; Zare et al., 2024). A simple example is behavior cloning (BC), which utilizes supervised learning to map observed situations to expert choices (Pomerleau, 1991). Building on the principles of imitation learning, the Decision Transformer (Chen et al., 2021) is similar to behavior cloning, framing reinforcement learning as sequence modeling, where an autoregressive Transformer is trained to predict the next action given a sequence rollout of returns, states, and actions. This view connects control to next-token prediction and attains competitive performance without explicit value function learning. The application of imitation learning extends beyond traditional domains. For instance, beyond robotics and games, IL has guided decision-making in exact optimization solvers. In mixed-integer linear programming, policies learned to imitate strong branching can be used within branch-and-bound and achieve strong results (Gasse et al., 2019). Related work also learns branching policies within branch-and-bound (B&B) for mixed-integer programming (MIP) (Zarpellon et al., 2021; He et al., 2014; Khalil et al., 2016). While these methods are close in spirit, the underlying search dynamics differ: MIP B&B explores a monotone tree with strong-branching oracles and immediate, node-local feedback, whereas SAT CDCL performs non-monotone search with restarts, clause learning, and delayed, non-local feedback. From the imitation-learning perspective, prior MIP approaches typically learn local ranking models on MIP features, whereas we train a transformer-based autoregressive policy that imitates entire CDCL KeyTraces. Thus, our setting complements IL for MIP branching, differing both in the target solver dynamics and in the learning formulation.
@@ 3 Preliminaries on SAT
and a formula is in conjunctive normal form (CNF) (Plaisted and Greenbaum, 1986) if it is a conjunction of clauses:
@@
In practice, SAT formulas are often represented in the standard DIMACS CNF format (Johnson and Trick, 1996). In this encoding, each clause is written as a sequence of nonzero integers terminated by 0, where the integer $i$ denotes variable $x_{i}$ and i-i denotes its negation $\lnot x_{i}$. For the formula in Equation 3, the DIMACS form is:
@@
Conflict-driven clause learning (CDCL) is the dominant algorithmic framework for practical SAT solvers (Silva and Sakallah, 1996; Biere et al., 2009). A CDCL solver incrementally explores a branching search tree through three fundamental operations: decision ($\mathtt{D}$), where the solver assigns a value to a chosen literal and thus extends the current partial assignment; unit propagation ($\mathtt{A}$), also known as Boolean Constraint Propagation (BCP) (Moskewicz et al., 2001), where implied literals are deduced from unit clauses; and backtracking ($\mathtt{BT}$), where a detected conflict triggers clause learning and a non-chronological jump to an earlier decision level.
@@
Usually, the choice of the next decision variable, together with its phase, is determined by a branching heuristic. The most influential family of heuristics is the Variable State Independent Decaying Sum (VSIDS) (Moskewicz et al., 2001). In VSIDS, variables are assigned activity scores that are increased whenever they appear in learned clauses. These scores are then periodically decayed, ensuring that recently relevant variables are favored for branching. This dynamic prioritization enables the solver to focus on the most promising parts of the search space. A notable implementation is MiniSAT (Eén and Sörensson, 2003), a lightweight and extensible open-source CDCL solver. Despite its minimalistic codebase, MiniSAT has become a standard platform for both research and industrial applications due to its clarity and effectiveness.
@@ 4.1 Expert KeyTrace Replay
We seek an expert that guides the solver toward effective branching decisions while avoiding wasted detours. Raw CDCL trails are often long and contain many decisions that are later undone by backtracking. An ideal expert should preserve only the decisions that survive and discard detours, thereby yielding clean training targets and eliminating steps that do not advance the search. In CDCL, unit propagation dominates runtime, typically accounting for 80%–90% of the total solving time (Zhang and Malik, 2002; Davis et al., 2008; Moskewicz et al., 2001), a trend we also confirm in Figure 2. This observation suggests that an expert which avoids wasted branches can substantially reduce propagation and thereby improve overall efficiency.
@@ 5 Experiments
In this section, we evaluate ImitSAT as a learned branching policy for CDCL solvers. We first describe the experimental setup, including baselines, solver implementation, datasets, model, and evaluation metrics. We then compare ImitSAT with SATformer and Graph-Q-SAT on random 3-SAT test sets and on a range of structured SAT families. Wall clock behaviour is studied both in Section 5.4 and through a direct comparison with pure MiniSAT in Appendix O, and we also examine generalization to industrial benchmarks from SATCOMP in Appendix M. Finally, we summarize several additional analyses whose full results are deferred to the appendices, including a query budget ablation, a GNN augmented variant, a Top-K masking versus fallback study, and integrations with more advanced solvers such as CaDiCaL (Biere et al., 2024a) and Kissat (Biere et al., 2024b), detailed in Appendices N, P, Q, and R.
@@ 5.1 Implementation Details
We compare against SATformer (Shi et al., 2023) and Graph‑Q‑SAT (Kurin et al., 2020), which are trained using their public implementations and specified training data. We also retrain them on our dataset, with equivalent training volume, denoted as SATformer* and GQSAT*.
@@
A Python reimplementation of MiniSAT 2.2 (Eén and Sörensson, 2003) is used and validated against the official C++ version on full trails $\mathcal{T}_{t}$. This version streamlines the integration of learning methods, ensuring that all approaches are evaluated in a consistent environment.
@@
We generate the random 3–SAT training dataset with a planted assignment (Mezard and Montanari, 2009; Achlioptas et al., 2000; 2005). The clause–variable ratio is in $[4.1,4.4]$. Variable ranges include 5 to 15, 16 to 30, 31 to 60, and 61 to 100, with additional fixed sizes of 50 and 100. For each formula, a MiniSAT run yields the level-annotated trail as Equation 5. Then, the expert KeyTrace is extracted using Equation 7. Every decision position provides one supervision pair, a KeyTrace prefix $\mathcal{K}_{t}$, and its next decision $d_{t+1}$. For evaluation, we first use the test set, which is drawn from the same generators and variable ranges as those used for training. Generalization is then assessed on different SAT families from SATLIB (Hoos and Stützle, 2000). Further dataset details appear in Appendix B.
@@
Propagation dominates CDCL time (Zhang and Malik, 2002; Davis et al., 2008; Moskewicz et al., 2001), as seen in Figure 2 and Table 6. We therefore use the number of propagation as the primary signal. Since instances vary in size and difficulty, we normalize by MiniSAT on a per‑instance basis and aggregate with the median to reduce the influence of outliers.
@@ 5.2 Comparison on test sets
We compare ImitSAT with Graph-Q-SAT and SATformer on held-out random 3-SAT. To align wall-clock budgets, ImitSAT is queried at 3 or 5 times per instance. Graph-Q-SAT uses the same computational budgets. SATformer updates the VSIDS variable scores once at initialization. We report MRPP $\tilde{r}$ and the one percent win rate $W_{1\%}$. We also include the GQSAT* and SATformer* variants, which are trained on our random 3-SAT dataset to have a fair comparison. On the test set, GQSAT* performs much worse than GQSAT, while SATformer* remains close to SATformer.
@@
MRPP $\tilde{r}$ ($\downarrow$) KeyTrace 0.57 0.39 0.21 0.10 0.17 0.03
GQSAT-3calls 1.00 0.94 0.89 1.15 0.71 0.85
@@
$W_{1\%}$ ($\uparrow$) KeyTrace 0.71 0.87 0.96 0.98 0.97 0.99
GQSAT-3calls 0.45 0.53 0.54 0.48 0.57 0.55
@@ 5.3 Generalization on special SAT families
Across all families, ImitSAT attains the lowest MRPP $\tilde{r}$ or ties for best as shown in Table 2 (top). The gains are large on PARITY and PRET, and with 5 calls ImitSAT is best on JNH and AIM, and matches the best on PHOLE. Graph-Q-SAT does not reduce propagation number on JNH and AIM, and in fact increases the propagation number. SATformer only shows a gain on the PARITY dataset; all other datasets exhibit no gain or worse performance than MiniSAT. Table 2 (bottom) reports the $W_{1\%}$. ImitSAT achieves the highest or tied win rate on all families, including a perfect score on PRET and matching the best on PARITY and PHOLE. These outcomes indicate that imitation of the expert KeyTrace transfers from random 3-SAT to structured regimes without finetuning. We report additional experiments on industrial benchmarks from SATCOMP (Iser and Jabs, 2024) in Appendix M, using instances with at most 100 variables and a DIMACS encoding that fits within our context budget. On this filtered subset, ImitSAT and SATformer solve the harder instances substantially faster in wall-clock time than Graph-Q-SAT under the same query budget as shwon in Figure 6.
@@
MRPP $\tilde{r}$ ($\downarrow$) KeyTrace 0.18 0.55 0.11 0.97 0.56
GQSAT-3calls 1.29 1.20 0.82 1.03 0.88
@@
$W_{1\%}$ ($\uparrow$) KeyTrace 1.00 0.75 1.00 0.50 1.00
GQSAT-3calls 0.38 0.31 0.80 0.50 0.50
@@ 5.4 Wall-clock Time
Across random 3‑SAT test sets and structured families, ImitSAT traces the lowest curves under these budgets, which means more instances are solved in less time, as shown in Figure 4. Learning model-based branching introduces query overhead, so wall‑clock gains appear only once propagation savings exceed this cost. We also compare ImitSAT directly with pure MiniSAT on 16–30, 31–60, and 61–100, as summarized in Table 12 and detailed in Appendix O. On the easier 16–30 and 31–60 ranges, MiniSAT remains faster while ImitSAT is the strongest learned method and stays close in runtime; on the harder 61–100 range, ImitSAT achieves the lowest wall‑clock time overall, showing that its propagation savings translate into a net speedup once instances are sufficiently large and challenging.
Figure 4: Wall‑clock time versus instances solved on random test sets and structured families. Curves show end‑to‑end solver time per instance with all model calls included and preprocessing excluded. ImitSAT and Graph‑Q‑SAT use 3 or 5 calls per instance. SATformer performs a single VSIDS initialization. The lower curve is better.
@@ 6 Conclusion
Acknowledgments
X. Xu acknowledges support from the National Natural Science Foundation of China (62302385).
@@ Appendix B Datasets
This section details the training corpus, held‑out test sets, and the structured SAT families used in evaluation. For the synthetic data, both train and test, all instances are random 3-SAT with a planted assignment and a clause–variable ratio in [4.1, 4.4][4.1,\,4.4] (Mezard and Montanari, 2009; Achlioptas et al., 2000; 2005). We categorize instances by the number of variables $n_{v}$ to control difficulty and balance the number of decision probes contributed by each category. Since larger $n_{v}$ produces longer trails and thus more supervision per instance, we allocate fewer instances at larger $n_{v}$. This ensures each bucket contributes a similar amount of training signal. We define buckets as ranges of variable counts (e.g., “5–15”), with each bucket grouping instances by the number of variables $n_{v}$ they contain.
@@
To assess cross‑distribution generalization, we evaluate on classic structured benchmarks from SATLIB (Hoos and Stützle, 2000). The suite spans both SAT and UNSAT, and includes 3-SAT and non–kk–SAT regimes. Specifically, we use AIM (Asahiro et al., 1996), which is equivalent to 3-SAT, focusing on the $n_{v}=100$ SAT portion. We also consider the SAT subset of JNH (Selman et al., 1996) with $n_{v}=100$ and PARITY (Warners and Van Maaren, 1998; Hoos and Stützle, 2000) in the compressed series, both as structured SAT families distinct from our random generator. Additionally, PRET (Johnson and Trick, 1996; Hoos and Stützle, 2000) is a two-coloring UNSAT problem with $n_{v}=60$, and the pigeonhole principle (PHOLE) (Haken, 1985) is used as a non–kk–SAT and UNSAT stress test. All subsets described above are considered with $n_{v}\leq 100$. Table 5 lists the families.